# On the self-consistency of compact objects in Lorentz-violating gravity theories

### v1

@@ Abstract
$ds^{2}=-A(r)dt^{2}+\frac{dr^{2}}{B(r)}+r^{2}d\Omega^{2}$ (10)
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the line element of a unit sphere. If the spacetime has horizons, they are found at the zeroes of $B(r)$—for instance in black hole spacetimes. On the other hand, the zeroes of $A(r)$ are surfaces with infinite redshift. For convenience, let us consider that $B(r)=A(r)/\Omega^{2}(r)$, where $\Omega^{2}(r)$ is a positive function of $r$. We assume that the above line element satisfies the corresponding Einstein field equations of the Lorentz-violating model (1) with some matter distribution (including vacuum) that does not couple to the Lorentz-violating fields.
Among the plenty of compact objects, such as black holes and stars that can be described by the above line element, it can also represent objects endowed with a minimal hypersurface, known as wormholes. The area of the minimal hypersurface is given by $S=4\pi a^{2}$, where $a$ is the so-called throat radius. Static and spherically symmetric wormholes can be conveniently described by the Morris-Thorne line element, which is obtained by performing $\Omega^{2}(r)=A(r)/(1-s(r)/r)$, where $s(r)$ is the so-called shape function. Additionally, it is convenient to cast the redshift function as $A(r)=\exp(2\Phi(r))$. In order to describe a singularity-free and traversable wormhole, the metric function must satisfy some additional requirements. To prevent the emergence of singularities, $\Phi(r)$ must be finite everywhere, in particular when $\Phi(r)=0$ the wormhole is said to be tideless. Additionally, the throat of the wormhole is located at the minimum of the radial coordinate, $r=a$, with the shape function satisfying $s(a)=a$. Moreover, $s(r)/r<1$ for $r>a$ provides the finiteness of the proper radial distance [^1] The proper radial distance is implicitly defined by $dx/dr=\pm[1-s(r)/r]^{-1/2}$, where $\pm$ refers to the two regions connected by the wormhole throat., $x$, throughout the spacetime.
Bumblebee constraints. First, let us consider the Einstein-bumblebee model (only $\tilde{\xi}_{2}\neq 0$) and that the dynamics of the bumblebee field is frozen in its VEV state, $\langle\mathfrak{B}_{\mu}\rangle=\mathfrak{b}_{\mu}$, and additionally consider a spacelike bumblebee configurations given by
@@
such configuration was considered, for instance, in Refs. [10, 13, 16]. This configuration has vanishing field strength, i.e., $\mathfrak{b}_{\mu\nu}=0$, and therefore the equation of the bumblebee field, in the VEV state, turns into a geometrical constraint:
$2\langle\tilde{V}_{Y}\rangle\mathfrak{b}^{\nu}-\frac{\tilde{\xi}_{2}}{\kappa}\mathfrak{b}_{\mu}R^{\mu\nu}=0,$ (12)
@@
$\Omega^{\prime}\left(rA^{\prime}+4A\right)-\Omega\left(rA^{\prime\prime}+2A^{\prime}\right)-\frac{4\kappa\langle\tilde{V}_{Y}\rangle r\Omega^{3}}{\tilde{\xi}_{2}}=0.$ (13)
@@
In particular, if $\Omega^{2}(r)$ is a constant, say $\Omega^{2}(r)=\omega^{2}$, the above constraint becomes
$-\omega\left(rA^{\prime\prime}+2A^{\prime}\right)-\frac{4\kappa\langle\tilde{V}_{Y}\rangle r\omega^{3}}{\tilde{\xi}_{2}}=0.$ (14)
If the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state, for instance a quadratic one, $\tilde{V}(Y)=\tfrac{1}{2}\tilde{\zeta}Y^{2}$, both $\langle\tilde{V}\rangle=0$ and $\langle\tilde{V}_{Y}\rangle=0$. Therefore, the above equation reduces to $rA^{\prime\prime}+2A^{\prime}=0$. Hence, to satisfy the constraint imposed by the bumblebee field in the considered VEV state, the redshift function must conform to
@@
where $a_{1,2}$ are integration constants that cannot be fixed solely from the constraint (12). In this scenario, only compact objects with a Schwarzschild-like redshift function are allowed by the anisotropy induced by the background bumblebee field. Particularly, one can find self-consistent black holes within the Einstein-bumblebee system. The static and spherically symmetric vacuum black hole solutions of the Einstein-bumblebee system obtained in Ref. [10] trivially satisfies the constraint (12). This black hole is described by $A(r)=1-2M/r$ and $B=A(r)/(1+l_{V})$, where $M$ is interpreted as the mass of the compact object and $l_{V}=\mathfrak{b}^{2}\tilde{\xi}_{2}$ contains the effects of the Lorentz violation.
On the other hand, if the potential is linear, say $\tilde{V}(Y)=\tfrac{1}{2}\tilde{\lambda}Y$, in the VEV state $\langle\tilde{V}\rangle$ vanishes but $\langle\tilde{V}_{Y}\rangle=\tfrac{1}{2}\lambda$. Therefore, the above equation reduces to $\tilde{\xi}_{2}(rA^{\prime\prime}+2A^{\prime})+2\kappa\omega^{2}\lambda r=0$. Consequently, to satisfy the constraint imposed by the bumblebee field in the considered VEV state, the redshift function must conform to
$A(r)=\tilde{a}_{1}-\frac{\tilde{a}_{2}}{r}-\frac{\omega^{2}\Lambda_{e}r^{2}}{3},$ (16)
that is the redshift function must have a Kottler-like form, where $\tilde{a}_{1,2}$ are integration constants that cannot be fixed solely by the constraint (12) and $\Lambda_{e}=\lambda\kappa/\tilde{\xi}_{2}$. As obtained in Ref. [13], under the same assumptions considered here, static and spherically symmetric vacuum solutions of the Einstein-bumblebee system with cosmological constant trivially satisfies the constraint (12), since the redshift function acquires the form (16), namely $A(r)=1-2M/r-(1+l_{V})\Lambda_{e}r^{2}/3$, and $B(r)=A(r)/(1+l_{V})$, where $M$ is interpreted as the mass of the compact object and $\Lambda_{e}$ an effective cosmological constant. As a consequence, black holes immersed in a universe with non-zero cosmological constant, are self-consistent objects within the Einstein-bumblebee system.
@@
$2\tilde{\xi}_{2}r^{2}[r-s(r)][\Phi^{\prime}(r)^{2}+\Phi^{\prime\prime}(r)]+4r^{3}\kappa\langle\tilde{V}_{Y}\rangle$
$+\tilde{\xi}_{2}[s(r)-rs^{\prime}(r))][2+r\Phi^{\prime}(r)]=0,$ (17)
@@
where we have rewritten the Eq. (13) in term of the Morris-Thorne metric functions, $\Phi(r)$ and $s(r)$. For simplicity, let us probe if tideless configurations, $\Phi(r)=0$, are self consistent in the Einstein-bumblebee model. In this case the above equation reduces to $\tilde{{\xi}_{2}}[s(r)-rs^{\prime}(r)]=-2r^{3}\kappa\langle\tilde{V}_{Y}\rangle$. It follows that, if the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state, the constraint is fulfilled only if $s(r)=s_{1}r$, where $s_{1}$ is a constant. Therefore $s(r)/r$ is constant regardless the value of $r$. It follows that, only if $s_{1}=1$ the requirement $s(a)/a=1$ holds, however this implies that $01-s(r)/r=$ 0 throughout the spacetime. If, on the contrary, $s_{1}\neq 1$ in order to satisfy $s(r)/r<1$ throughout the spacetime, $s(a)/a\neq 1$. This implies that the areal radius $r(x)$ does not have a local minimum, and therefore there is no throat in the spacetime. Consequently, if the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state and the bumblebee field is frozen at the considered VEV configuration, tideless wormholes are not self-consistent objects in the Einstein-bumblebee model. This shows that the tideless wormhole reported in Ref. [16], although it satisfies the corresponding Einstein equations, is not self-consistent within the Einstein-bumblebee model and cannot be cast as a solution of this model.
Let us instead consider that the potential that drives the Lorentz symmetry breaking is linear, for instance $\tilde{V}(Y)=\tfrac{1}{2}\lambda Y$. In this case, the constraint is fulfilled only if $s(r)=\tilde{s}_{1}r+r^{3}\Lambda_{e}/2$, where $\tilde{s}_{1}$ is a constant and $\Lambda_{e}$ is defined above. One can check that the condition $s(a)/a=1$ holds for $\tilde{s}_{1}=1-a^{2}\Lambda_{e}/2$. Additionally, if $\Lambda_{e}<0$, the condition $01-s(r)/r>$ 0 for $r>a$ also holds. For the sake of completeness, one can compute the proper radial distance for the allowed shape function, namely $x=\pm\pi/\sqrt{2\Lambda_{e}}\pm\sqrt{2/|\Lambda_{e}|}\arctan(r/\sqrt{r^{2}-a^{2}})$, where we have set $x(a)=0$. Therefore, the square of the areal radius, in terms of the proper radial distance, reads $r^{2}(x)=a^{2}\cosh^{2}(x\sqrt{|\Lambda_{e}|/2})$, and, as one can see, has a local minimum at $x=0$. Hence, considering the linear potential, one can find self-consistent tideless wormhole configurations within the Einstein-bumblebee model. It is important to point out that such allowed wormhole configurations in the Einstein-bumblebee model do not correspond to Ellis-Bronnikov-like wormholes. This is because the areal radius of the latter relates to its proper radial distance through a parabolic equation, specifically $r^{2}(x)=a^{2}+r_{1}x^{2}$, where the constant $r_{1}=1$ in the massless Ellis-Bronnikov wormhole.

### v2

@@ III Compact objects in Lorentz-violating gravity
$ds^{2}=-A(r)dt^{2}+\frac{dr^{2}}{B(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (9)
If $A(r)=0$, there is a surface in the spacetime marking the static limit: within regions where $A(r)<0$, no static observers can exist. The surface defined by $B(r)=0$ is a candidate horizon surface, as it implies the coordinate velocity $dr/dt$ for radial null geodesics vanishes. However, this condition alone is insufficient to define a true event horizon. An event horizon must additionally be a null hypersurface and form a causal boundary—verified by the global spacetime structure—such that no future-directed causal curves from within can reach future null infinity. In standard spherically symmetric black hole scenarios, for instance Schwarzschild or Reissner-Nordström black holes (where $A(r)=B(r)$), the locus of $B(r)=0$ coincide with those of $A(r)=0$, indicating that it represents a true event horizon. We remark that, if $B(r)=0$ occurs where $A(r)\neq 0$, the surface represents a minimal area hypersurface, not a horizon. For convenience, let us consider that $B(r)=A(r)/\Omega^{2}(r)$, where $\Omega^{2}(r)$ is a positive function of $r$. We assume that the above line element satisfies the corresponding Einstein field equations of the Lorentz-violating model (1) with some matter distribution (including vacuum) that does not couple to the Lorentz-violating fields.
Traversable wormholes represent a special class of horizonless solutions featuring a minimal surface (throat). These can conveniently be described by the Morris-Thorne line element [54],namely
$ds^{2}=-e^{2\Phi(r)}dt^{2}+\frac{dr^{2}}{1-s(r)/r}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (10)
which is obtained by casting the redshift function as $A(r)=\exp(2\Phi(r))$ and performing $\Omega^{2}(r)=A(r)/(1-s(r)/r)$, where $s(r)$ is the shape function. In order to describe a singularity-free and traversable wormhole, the metric function must satisfy some additional requirements [54, 55]. For instance, to prevent the emergence of singularities and horizons, $\Phi(r)$ must be finite everywhere. In particular, when $\Phi(r)=0$ the wormhole is said to be tideless. Additionally, the throat of the wormhole is located at the minimum of the radial coordinate, $r=a$, with the shape function satisfying
$s(a)=a,$ (11)
$s(r)/r<1\quad\text{for}\quad r>a.$ (12)
The first condition ensures $r=a$ is the minimum radius, while the second condition guarantees the finiteness of the proper radial distance $x$. This distance is implicitly defined by
$\frac{dx}{dr}=\pm\left[1-\frac{s(r)}{r}\right]^{-1/2},$ (13)
where $\pm$ refers to the two distinct spatial regions connected by the wormhole throat.
IV Bumblebee constraints
First, let us consider the Einstein-bumblebee model ($\xi_{1,2,3}=0$) and that the dynamics of the bumblebee field is frozen in its VEV state, $\langle\mathfrak{B}_{\mu}\rangle=\mathfrak{b}_{\mu}$, and additionally consider a spacelike bumblebee configurations given by
@@ IV Bumblebee constraints
such configuration was considered, for instance, in Refs. [30, 33, 50, 51]. This configuration has vanishing field strength, i.e., $\mathfrak{b}_{\mu\nu}=0$, and therefore the equation of the bumblebee field, in the VEV state, turns into a geometrical constraint:
$2\langle\tilde{V}_{Y}\rangle\mathfrak{b}^{\nu}-\frac{\tilde{\xi}_{1}}{\kappa}R\mathfrak{b}^{\nu}-\frac{\tilde{\xi}_{2}}{\kappa}\mathfrak{b}_{\mu}R^{\mu\nu}=0,$ (15)
@@
$\frac{\tilde{\xi}_{1}}{r}[2r^{2}A^{\prime}\Omega^{\prime}-2r\Omega\left(rA^{\prime\prime}+4A^{\prime}\right)-4A\left(\Omega-2r\Omega^{\prime}\right)+4\Omega^{3}]$
$+\tilde{\xi}_{2}[\Omega^{\prime}\left(rA^{\prime}+4A\right)-\Omega\left(rA^{\prime\prime}+2A^{\prime}\right)]-4\kappa\langle\tilde{V}_{Y}\rangle r\Omega^{3}=0.$ (16)
@@
We first address the case where $\Omega^{2}(r)$ is constant, specifically $\Omega^{2}(r)=\omega^{2}$, which transforms the aforementioned constraint into
$\tilde{\xi}_{1}\bigg{[}2rA^{\prime\prime}+\frac{4\left(2rA^{\prime}+A-\omega^{2}\right)}{r}\bigg{]}+\tilde{\xi}_{2}(rA^{\prime\prime}+2A^{\prime})$
$+4\kappa\langle\tilde{V}_{Y}\rangle r\omega^{2}=0.$ (17)
In particular, let us first consider the most studied Einstein-bumblebee model, where $\tilde{\xi}_{1}=0$. If the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state, for instance a quadratic one, $\tilde{V}(Y)=\tfrac{1}{2}\tilde{\zeta}Y^{2}$, both $\langle\tilde{V}\rangle=0$ and $\langle\tilde{V}_{Y}\rangle=0$. Therefore, the above equation reduces to $rA^{\prime\prime}+2A^{\prime}=0$. Hence, to satisfy the constraint imposed by the bumblebee field in the considered VEV state, the redshift function must conform to
@@
where $a_{1,2}$ are integration constants that cannot be fixed solely from the constraint (15). In this scenario, only compact objects with a Schwarzschild-like redshift function are allowed by the the background bumblebee field. Particularly, one can find self-consistent black holes within the Einstein-bumblebee system. The static and spherically symmetric vacuum black hole solutions of the Einstein-bumblebee system obtained in Ref. [30] trivially satisfies the constraint (15). This black hole is described by A(r)=12M/rA(r)=1-2M/r and $B=A(r)/(1+l_{V})$, where $M$ is interpreted as the mass of the compact object and $l_{V}=\mathfrak{b}^{2}\tilde{\xi}_{2}$ contains the effects of the Lorentz violation. As reported in Ref. [56], this deviation from the spontaneous Lorentz symmetry breaking is connected to a geometrical effect of a conical angle $\Delta=\pi^{2}/\sqrt{1+l_{V}}$. Asymptotically, these black holes do not approach a flat region, rather their equatorial plane display a conical geometry. This produces observable effects like deviations in light deflection angles, even in the weak-field regime [30].
On the other hand, if the potential is linear, say $\tilde{V}(Y)=\tfrac{1}{2}\tilde{\lambda}Y$, in the VEV state $\langle\tilde{V}\rangle$ vanishes but $\langle\tilde{V}_{Y}\rangle=\tfrac{1}{2}\tilde{\lambda}$. Therefore, the above equation reduces to $\tilde{\xi}_{2}(rA^{\prime\prime}+2A^{\prime})+2\kappa\omega^{2}\tilde{\lambda}r=0$. Consequently, to satisfy the constraint imposed by the bumblebee field in the considered VEV state, the redshift function must conform to
$A(r)=\tilde{a}_{1}-\frac{\tilde{a}_{2}}{r}-\frac{\omega^{2}\tilde{\Lambda}_{e}r^{2}}{3},$ (19)
that is the redshift function must have a Kottler-like form, where $\tilde{a}_{1,2}$ are integration constants that cannot be fixed solely by the constraint (15) and $\tilde{\Lambda}_{e}=\tilde{\lambda}\kappa/\tilde{\xi}_{2}$. As obtained in Ref. [33], under the same assumptions considered here, static and spherically symmetric vacuum solutions of the Einstein-bumblebee system with cosmological constant trivially satisfies the constraint (15), since the redshift function acquires the form (19), namely $A(r)=1-2M/r-(1+l_{V})\tilde{\Lambda}_{e}r^{2}/3$, and $B(r)=A(r)/(1+l_{V})$, where $M$ is interpreted as the mass of the compact object and $\tilde{\Lambda}_{e}$ an effective cosmological constant. As a consequence, black holes immersed in a universe with non-zero cosmological constant, are self-consistent objects within the Einstein-bumblebee system.
For concreteness, let us also consider the less investigated Einstein-bumblebee model with vanishing $\tilde{\xi}_{2}$ (rather than vanishing $\tilde{\xi}_{1}$). In this case, the constraint (IV) reduces to
$2rA^{\prime\prime}+\frac{4(2rA^{\prime}+A-\omega^{2})}{r}=-\frac{4\kappa\langle\tilde{V}_{Y}\rangle r\omega^{2}}{\tilde{\xi}_{1}}.$ (20)
If the potential driving Lorentz symmetry breaking extremizes at the vacuum state, the lapse function must conform to $A(r)=\omega^{2}-\alpha_{1}/r-\alpha_{2}/r^{2}$, where $\alpha_{1,2}$ are constants. Conversely, for the case of a linear potential ($\tilde{V}=\tfrac{1}{2}\tilde{\lambda}Y$), the lapse function must conform to $A(r)=\omega^{2}-\alpha_{2}/r-\alpha_{3}/r^{2}+\kappa\tilde{\lambda}\omega^{2}r^{2}/12$. To the best of our knowledge, compact object solutions with such lapse function structures have not yet been reported in the literature, motivating further investigation in this direction. We point out that, for simultaneous non-zero $\tilde{\xi}_{1}$ and $\tilde{\xi}_{2}$, analytical investigations of the constraint (IV) are challenging.
@@
$4\tilde{\xi}_{1}r^{2}(r-s)\left(\Phi^{\prime\prime}+\Phi^{\prime 2}\right)-2\tilde{\xi}_{1}r[\left(r\left(s^{\prime}-4\right)+3s\right)\Phi^{\prime}+2s^{\prime}]$
$+2\tilde{\xi}_{2}r^{2}[r-s][\Phi^{\prime 2}+\Phi^{\prime\prime}]+\tilde{\xi}_{2}[s-rs^{\prime}][2+r\Phi^{\prime}]$
$+4r^{3}\kappa\langle\tilde{V}_{Y}\rangle=0,$ (21)
@@
where we have rewritten the Eq. (IV) in term of the Morris-Thorne metric functions, $\Phi(r)$ and $s(r)$. For simplicity, let us probe if tideless configurations, $\Phi(r)=0$, are self consistent within the Einstein-bumblebee model. In this case the above equation reduces to $2\tilde{\xi}_{1}rs^{\prime}+\tilde{{\xi}}_{2}[s-rs^{\prime}]=-2r^{3}\kappa\langle\tilde{V}_{Y}\rangle$. It follows that, if the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state, the constraint is fulfilled only if $s(r)=s_{1}r^{\beta}$, where $s_{1}$ is a constant and $\beta=\tilde{\xi}_{2}/(\tilde{\xi}_{2}+2\tilde{\xi}_{1})$. In order to satisfy $s(r)/r<1$ throughout the spacetime and $s(a)/a=1$, it follows that $s_{1}=a^{1-\beta}$ and $\beta<1$. Remarkably, if the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state and the bumblebee field is frozen at the considered VEV configuration, tideless wormholes are not self-consistent objects in the Einstein-bumblebee model with $\tilde{\xi}_{1}=0$. Since $\beta=1$ ($\tilde{\xi}_{1}=0$), $s(r)/r$ is constant regardless the value of $r$ and therefore there is no throat in the spacetime.
The above result can be summarized as follows: Within Einstein-bumblebee gravity incorporating non-minimal couplings to the Ricci tensor—described by the action
$S$ $=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa}(R-2\Lambda)-\frac{1}{4}\mathfrak{B}_{\mu\nu}\mathfrak{B}^{\mu\nu}-\tilde{V}(Y)\right$.
$\left.+\frac{\tilde{\xi}_{2}}{2\kappa}\mathfrak{B}^{\mu}\mathfrak{B}^{\nu}R_{\mu\nu}\right]+S_{m},$ (22)
—the spontaneous local Lorentz and diffeomorphism symmetry breaking of the bumblebee field, manifesting as a non-zero VEV specified in Eq. (14), prohibits the existence of tideless wormholes when the bumblebee self-interaction potential extremizes at this VEV configuration. This is particularly relevant because this model represents one of the most investigated gravitational frameworks exhibiting spontaneous local Lorentz and diffeomorphism symmetry breaking. Remarkably, this shows that the tideless wormhole reported in Ref. [50], although it satisfies the corresponding Einstein equations, is not self-consistent within the considered Einstein-bumblebee model and cannot be cast as a solution of this model.
We emphasize that, provided $\tilde{V}$ extremizes at the VEV, self-consistent tideless wormholes can exist when $\tilde{\xi}_{1}\neq 0$. Two distinct classes emerge, depending on the specific bumblebee framework, namely
  • For $\tilde{\xi}_{1}\neq 0$ and $\tilde{\xi}_{2}\neq 0$: $s(r)=a^{1-\beta}r^{\beta}$ with $\beta\equiv\dfrac{\tilde{\xi}_{2}}{\tilde{\xi}_{2}+2\tilde{\xi}_{1}}<1$;
  • For $\tilde{\xi}_{1}\neq 0$ and $\tilde{\xi}_{2}=0$: $s(r)=a$ (constant shape function).
Conversely, when considering that the potential that drives the Lorentz symmetry breaking is linear ($\tilde{V}(Y)=\tfrac{1}{2}\tilde{\lambda}Y$), the constraint is fulfilled only if $s(r)=s_{2}r^{\beta}+r^{3}\kappa\tilde{\lambda}/(6\tilde{\xi}_{1}+2\tilde{\xi}_{2})$. To $s(r)/r<1$ throughout the spacetime and $s(a)/a=1$, one requires that $s_{2}=a^{1-\beta}[1-a^{2}\kappa\tilde{\lambda}/(6\tilde{\xi}_{1}+2\tilde{\xi}_{2})]$, with $\kappa\tilde{\lambda}/(6\tilde{\xi}_{1}+2\tilde{\xi}_{2})<0$ and $\beta\leq 1$. Hence, differently from the case where $\tilde{V}$ extremizes at the VEV state, one can find self-consistent tideless wormhole configurations within the Einstein-bumblebee model with $\tilde{\xi}_{1}=0$ ($\beta=1$).
Before concluding this section, let us comment on the consistency of previously reported interior solutions for compact stars within the bumblebee model [51, 57]. While the exterior solution reported in Ref. [51, 57]—where the authors considered the bumblebee model with $\tilde{\xi}_{1}=0$ and $\langle\tilde{V}_{Y}\rangle=0$—matches the vacuum solution of Ref. [30] and, from our analysis, is consistent, the interior solution requires careful attention. To determine the interior solution, the authors solve the bumblebee-modified Tolman–Oppenheimer–Volkoff (TOV) equation without considering the constraint (IV). This can lead, as we have seen, to inconsistencies.
To illustrate this issue, consider the analytical constant-density compact star solution reported in Ref. [51] with interior solution
$A(r)=$ $\left[\frac{3}{2}\sqrt{1-\frac{2M_{\star}}{R_{\star}}}-\frac{1}{2}\sqrt{1-\frac{2M_{\star}r^{2}}{R_{\star}^{3}}}\right$.
$\left.+\frac{3l_{V}}{4}\left(\sqrt{1-\frac{2M_{\star}r^{2}}{R_{\star}^{3}}}-\sqrt{1-\frac{2M_{\star}}{R_{\star}}}\right)\right]^{2},$ (23)
$B(r)$ $=\frac{A(r)}{\Omega^{2}(r)}=\frac{1}{1+l_{V}}\left(1-2m_{0}r^{2}\right),$ (24)
where $M_{\star}$ and $R_{\star}$ denote the star’s mass and radius, and $m_{0}=4\pi\rho_{0}/3$, where $\rho_{0}$ is the constant density of the star. Substituting this $B(r)$ into constraint (IV) yields:
$A(-2A^{\prime\prime}+4m_{0}r(A^{\prime}+rA^{\prime\prime}))$
$+16m_{0}A^{2}+(1-2m_{0}r^{2})(A^{\prime})^{2}=0.$ (25)
The only redshift function $A(r)$ satisfying this constraint is
$A(r)=\beta_{1}\cosh^{2}\left[\sqrt{2}\arcsin(\sqrt{2m_{0}}r)+\beta_{2}\right],$ (26)
for constants $\beta_{1},\beta_{2}$. This differs fundamentally from the form in (23), demonstrating that the reported solution violates constraint (IV) and is thus inconsistent within the bumblebee model. Consequently, solving the modified TOV equations alone—without enforcing the Lorentz-violating field equations—is insufficient to establish valid compact star solutions in this theory.
We emphasize that there is a significant challenge in determining whether compact objects with non-constant $\Omega^{2}(r)$ satisfy the constraint (IV). This case generally requires numerical methods unless additional assumptions (like the tideless condition above) are imposed. However, as our analysis does not assume a specific matter distribution, numerical investigations focusing solely on the bumblebee field equations risk producing ambiguous or biased results. For this reason, we have restricted our study to analytical methods. Our analysis, nevertheless, stresses that any future numerical work must solve all field equations (including those generating geometric constraints) simultaneously to ensure solution consistency (see for instance the successful implementation in Ref. [34]).
V Antisymmetric rank-2 tensor constraints